# Text-to-Remote-Sensing-Image Retrieval beyond RGB Sources

### v1

@@ Abstract
Retrieving relevant imagery from vast satellite archives is crucial for applications like disaster response and long-term climate monitoring. However, most text-to-image retrieval systems are limited to RGB data, failing to exploit the unique physical information captured by other sensors, such as the all-weather structural sensitivity of Synthetic Aperture Radar (SAR) or the spectral signatures in optical multispectral data. To bridge this gap, we introduce CrisisLandMark, a new large-scale corpus of over 647,000 Sentinel-1 SAR and Sentinel-2 multispectral images paired with structured textual annotations for land cover, land use, and crisis events harmonized from authoritative land cover systems (CORINE and Dynamic World) and crisis-specific sources. We then present CLOSP (Contrastive Language Optical SAR Pretraining), a novel framework that uses text as a bridge to align unpaired optical and SAR images into a unified embedding space. Our experiments show that CLOSP achieves a new state-of-the-art, improving retrieval nDGC by 54% over existing models. Additionally, we find that the unified training strategy overcomes the inherent difficulty of interpreting SAR imagery by transferring rich semantic knowledge from the optical domain with indirect interaction. Furthermore, GeoCLOSP, which integrates geographic coordinates into our framework, creates a powerful trade-off between generality and specificity: while the CLOSP excels at general semantic tasks, the GeoCLOSP becomes a specialized expert for retrieving location-dependent crisis events and rare geographic features. This work highlights that the integration of diverse sensor data and geographic context is essential for unlocking the full potential of remote sensing archives.
@@
Text-to-Image Retrieval , Remote Sensing , Geospatial Data , Cross-Modal Retrieval , Crisis Management , Land Use and Land Cover
journal: Remote Sensing of Environment
@@
[polito]organization=Politecnico di Torino,addressline=Corso Duca Degli Abruzzi, city=Turin, postcode=10129, country=Italy
@@ 1 Introduction
Satellite remote sensing data have become paramount in several application domains, such as climate and environmental monitoring, disaster management, urban planning, and agriculture (Wulder et al., 2022; Torres et al., 2012; Drusch et al., 2012; Goward et al., 2001; Roy et al., 2014). However, due to the enormous scale and diversity of these data, searching for relevant content is challenging, especially when queries are expressed in natural language rather than by visual examples or geospatial coordinates. To tackle this issue, Text-to-Remote-Sensing-Image Retrieval (T2RSIR, in short) aims to leverage the textual and geospatial annotations associated with the visual content to retrieve the satellite images that are most relevant to a natural language query. Figure 1 depicts the main steps of the T2RSIR process, i.e., indexing, retrieval, and ranking.
While recent T2RSIR solutions mainly focus on RGB sources and high-resolution aerial imagery (Wang et al., 2024; Jain et al., 2024; Lu et al., 2018; Yuan et al., 2022; Qu et al., 2016), these approaches fail to exploit the richer data and global coverage available in satellite missions. This reliance on optical, often aerial, imagery limits our ability to perform large-scale analysis under diverse atmospheric conditions and overlooks crucial biophysical information hidden in the visible spectrum. Specifically, missions like Sentinel-1 provide SAR data, which are robust to cloud coverage and atmospheric conditions (Torres et al., 2012) and particularly suited to address tasks like land cover mapping (Clasen et al., 2024; Abdikan et al., 2016), agricultural applications (Sainte Fare Garnot and Landrieu, 2021; Turkoglu et al., 2021), change detection (Rege Cambrin and Garza, 2024), and infrastructure monitoring (Li et al., 2020). Instead, Sentinel-2 supplies multispectral bands (Drusch et al., 2012), which are crucial for applications like mapping vegetation (Thenkabail et al., 2004), water bodies (Gao, 1996), and soil characteristics (Huete, 1988). Unlike aerial imagery captured from aircraft at lower altitudes, which is high-resolution and has localized scope, satellite imagery covers global scales with consistent acquisition frequencies, making it indispensable for large-scale applications.
Unfortunately, public remote sensing datasets typically lack either (i) the full range of spectral or SAR information (beyond RGB) or (ii) sufficient textual annotations, limiting their usability for T2RSIR. To bridge the gap, we propose a new corpus for T2RSIR, i.e., CrisisLandMark, consisting of over 647,000 images acquired by Sentinel-1 (SAR) and Sentinel-2 (optical) spatial missions. The corpus is enriched with structured textual annotations derived from authoritative sources and crisis-event tags. Rather than relying on potentially ambiguous free-form text, we generated labels by harmonizing two established land cover classification systems: the detailed CORINE Land Cover (CLC) system (Büttner et al., 2017) for European regions and the global, near-real-time Dynamic World (DW) system (Brown et al., 2022). They provide salient information about the earth’s biophysical cover (land cover), the way humans utilize these covers (land use), and information for emergency management (crisis events), which are generally missing from many datasets. Importantly, structured annotations allow efficient analysis of land changes and a deeper exploration and explanation of the results produced by models trained on this dataset. Additionally, unlike the unstructured captions common in existing benchmarks (Wang et al., 2024; Lu et al., 2018; Yuan et al., 2022; Qu et al., 2016), our structured tags provide unambiguous, machine-readable semantic information. This mitigates ambiguity and allows a quantitatively verifiable interpretation of results.
To effectively and efficiently address T2RSIR on the new corpus, we train and release a contrastive multimodal learning architecture inspired by CLIP (Radford et al., 2021), which encodes the image-text pairs into a shared latent space. The newly proposed network, namely CLOSP (Contrastive Language Optical SAR Pretraining), aligns Sentinel-1 and Sentinel-2 images annotated with land use, land cover, and crisis events information, allowing for robust retrieval across diverse conditions and sensor types. CLOSP is the first multimodal retrieval architecture designed to create a shared semantic space between textual descriptions, optical multispectral imagery, and SAR data, explicitly moving beyond the limitations of RGB-only systems. Additionally, to overcome the challenge that SAR and optical images are rarely captured at the same time, CLOSP uses the shared text annotations as an anchor to bridge the two visual modalities. This strategy aligns each sensor’s data with the text independently, effectively creating a unified semantic space without requiring direct, time-matched image pairs. Furthermore, we introduce GeoCLOSP, an extension that integrates geographic coordinates. Our results reveal a distinct trade-off between semantic generality and geographic specificity: while the baseline CLOSP model excels at general-purpose semantic tasks, GeoCLOSP acts as a specialist model that improves performance on location-dependent queries, such as crisis events and rare geographic features.
@@
The experimental results on CrisisLandMark show that CLOSP consistently and substantially outperforms state-of-the-art T2RSIR baselines across multiple retrieval metrics. We demonstrate that our unified training strategy is highly effective, enabling knowledge transfer that significantly improves the semantic interpretation and retrieval of challenging SAR imagery. The achieved results underscore the value of integrating diverse sensor data like SAR and optical imagery for building more powerful and robust large-scale remote sensing retrieval systems.
The main paper contributions can be summarized as follows:
  • 1.
@@
    A new T2RSIR corpus (CrisisLandMark) including multispectral and radiometric information beyond RGB data with structured annotations for land use, land cover, and crisis events.
  • 2.
    A new multimodal learning architecture (CLOSP) that aligns (1) Text about land cover, land use, and crisis events, (2) Optical and SAR satellite imagery, and optionally (3) Geographical locations.
  • 3.
    An empirical validation on CrisisLandMark that (1) demonstrates the superiority of CLOSP over state-of-the-art models through knowledge transfer to the SAR domain, and (2) quantifies the trade-off between semantic and geographic relevance with GeoCLOSP.
The project code and the corpus are available for research purposes at https://github.com/DarthReca/closp.
2 Related Works
This section analyzes previous works on remote sensing, text-to-image retrieval, and their intersection.
2.1 T2RSIR Corpora
Due to the large amount of remote sensing imagery, many T2RSIR corpora were proposed to evaluate the effectiveness of retrieval models. However, many annotated corpora suffer from one or more shortcomings: (i) relatively small sizes (e.g.; RSICD (Lu et al., 2018), RSITMD (Yuan et al., 2022), UCM-Captions (Qu et al., 2016) have less than 11,000 samples), (ii) reliance on crowdsourcing or non-expert annotations (e.g.; SkyScript (Wang et al., 2024)), or (iii) captions generated with Large Language Models (Liu et al., 2025), which are subject to bias and hallucination effects (Caffagni et al., 2024). Additionally, they lack structured LULC annotations that mitigate the linguistic ambiguity inherent in free-form captions to provide a robust quantitative analysis. Note that, although Sen2Lucas (Jain et al., 2024) contains LULC annotations, it does not contain crisis-related annotations. Furthermore, all these corpora are limited to RGB imagery, discarding the richer spectral and SAR data available from modern satellites. Beyond T2RSIR corpora, many recently released datasets for tasks like classification and segmentation contain annotations for LULC (Maggiori et al., 2017; Shermeyer et al., 2020; Clasen et al., 2024) or crisis-related events (e.g.; CaBuAr (Cambrin et al., 2023), QuakeSet (Rege Cambrin and Garza, 2024), MMFlood (Montello et al., 2022), and Sen12Flood (Rambour et al., 2020)). However, they either lack textual annotations or are limited to a single phenomenon, and they are not suited to address the T2RSIR task without further refinements. To address these issues, we propose CrisisLandMark, a corpus of over 647,000 satellite images from Sentinel-1 (SAR) and Sentinel-2 (optical), paired with structured LULC and crisis-event annotations. It offers an integrated source of text, images, and locations that is one of a kind within the remote sensing community. Table 1 compares the different features of CrisisLandMark with existing T2RSIR benchmarks.
2.2 T2RSIR Models
Retrieving satellite imagery that semantically corresponds to a natural language query is a challenging task that has benefited from advances in multimodal learning. To solve this problem, the use of contrastive multimodal learning architectures like CLIP (Ilharco et al., 2021; Radford et al., 2021) is established. For example, SkyCLIP (Wang et al., 2024) extends the pretraining of the CLIP model using a subset of the SkyScript dataset to add in-domain knowledge. SenCLIP (Liu et al., 2024) aligns satellite RGB image representation to the ground-level visual knowledge of the pretrained CLIP model. RemoteCLIP (Jain et al., 2024) aligns satellite images with captions generated from object detection boxes and segmentation masks, but is limited to the RGB visual domain. SatCLIP (Klemmer et al., 2024) leverages multispectral satellite data, aligning them exclusively with the corresponding geographic coordinates. Unlike all prior models, CLOSP integrates both optical (Sentinel-2) and SAR (Sentinel-1) imagery, along with structured textual annotations covering land use, land cover, crisis events, and, optionally, geospatial coordinates. This design allows CLOSP to capture richer physical information from satellite sources and address broader application scenarios than prior methods.
@@ 3.1 Data sources
We enrich the source images with a comprehensive set of textual and geospatial annotations. To support crisis management applications, we incorporated the four specialized datasets (CaBuAr, MMFlood, Sen12Flood, Quakeset), retaining their original event tags (wildfire, flood, earthquake). For these crisis images, we generated corresponding LULC annotations by querying the Dynamic World collection. For general LULC coverage in re-BEN, we mapped the high-quality CORINE LC provided with the dataset as detailed in Section 3.3. Spatial coordinates also characterize all the image patches of every dataset.
The resulting corpus is primarily composed of general land cover scenes from re-BEN, which constitute 88% of the images, providing a robust foundation for learning diverse visual semantics. The remaining 12% from the specialized datasets offer critical examples for crisis retrieval tasks. Table 2 provides a detailed breakdown of the data sources and their respective contributions of Sentinel-1 and Sentinel-2 imagery.
@@ 3.2 Source Images
Sentinel-2 Level-2A (L2A) products (Drusch et al., 2012; European Space Agency, 2024b) are high-resolution satellite images that provide detailed, true-color views of the Earth’s surface, optimized for environmental and land monitoring. These images are processed to correct atmospheric effects like haze or cloud cover, making them more accurate and ready for analysis. They include 13 spectral bands, ranging from visible light (what we see) to infrared, enabling insights into vegetation health, water quality, soil conditions, and urban development. After the atmospheric correction, one band was removed, resulting in 12 bands, as seen in Table 3. One example image of the RGB bands can be seen in LABEL:fig:example_general_S2 paired with VV and VH channels (LABEL:fig:example_general_S1_1 and LABEL:fig:example_general_S1_2) of a Sentinel-1 GRD product of the same area. The different vision achieved with different spectral bands is shown in Figure 3 in the case of a wildfire.
Figure 2: Sample images from Sentinel-1 VV, and Sentinel-1 VH, and Sentinel-2 RGB of the same area. The scale of Sentinel-1 image values for each channel is the same.
@@ 3.3 Source text
Since CORINE is limited to specific years and only available for Europe, we enrich the four crisis-event datasets (that cover different years and areas than CORINE) with the 9 LULC classes of Dynamic World (DW) (Brown et al., 2022) (i.e., Water, Trees, Grass, Flooded Vegetation, Crops, Shrub and Scrub, Built area, Bare Ground, Snow), which are contained in CORINE categories. The crisis annotation is taken from the original dataset. Dynamic World is a high-resolution, near-real-time global land cover system developed by Google, leveraging deep learning models and Sentinel-2 satellite imagery. It provides 10-meter spatial resolution land cover classifications of the entire globe, making it applicable to diverse ecosystems and regions. Unlike traditional datasets, it offers frequent updates, enabling the detection of rapid and fine-scale land cover changes. The crisis datasets also contain special keywords to denote burned area, earthquake damaged, and flooded area. To create a unified and scalable label space, we harmonized all annotations using the 9-class system from DW. Thanks to its near-real-time global coverage, it is ideal for analyzing diverse ecosystems and detecting rapid land cover changes. With this solution, we map the higher quality of CLC annotation to the same classes of DW. The description and distribution of the classes can be found in Table 4. The mapping is detailed in A.
@@ 3.4 Queries
We define the query set by taking every unique combination (with length from 1 to 12) of one or more labels that co-occur in at least one image within our corpus. This process resulted in a total of 2,047 distinct multi-label queries. We define a graded relevance $rel$ based on the IoU between the query labels $L_{q}$ and the image labels $L_{i}$ to provide a more fine-grained evaluation:
@@
For binary relevance operators (e.g., precision and recall), we consider a threshold of 5 for defining a relevant item. The distribution of relevance scores and relevant images per query is shown in Figure 4.
@@ 4 Methodology
Figure 5: The CLOSP model aligns textual descriptions with SAR and optical satellite images. For a batch of $N$ elements ($M$ SAR and $M$ MSI), one modality—either SAR or optical—is selected for each element of the batch, and the corresponding image embeddings are paired with their associated textual embeddings. The model is trained to maximize alignment for these positive pairs (represented by white cells in the matrix) while ensuring that negative pairs, formed by combining textual and image embeddings from different items within the batch (represented by grey cells in the matrix), are effectively separated. GeoCLOSP extends the CLOSP architecture by incorporating a location encoder, which aligns the geographical coordinates of an item with the corresponding satellite image, in parallel with the image-text alignment.
The remainder of this section is organized as follows. Section 4.1 formulates the task, whereas Sections 4.2 and 4.3 respectively describe CLOSP and its variant that also leverages geospatial data (GeoCLOSP).

### v2

# Visualization, Software, Conceptualization, Writing - Original Draft

@@ Abstract
Retrieving relevant imagery from vast satellite archives is crucial for applications like disaster response and long-term climate monitoring. However, most text-to-image retrieval systems are limited to RGB data, failing to exploit the unique physical information captured by other sensors, such as the all-weather structural sensitivity of Synthetic Aperture Radar (SAR) or the spectral signatures in optical multispectral data. To bridge this gap, we introduce CrisisLandMark, a new large-scale corpus of over 647,000 Sentinel-1 SAR and Sentinel-2 multispectral images paired with structured textual annotations for land cover, land use, and crisis events harmonized from authoritative land cover systems (CORINE and Dynamic World) and crisis-specific sources. We then present CLOSP (Contrastive Language Optical SAR Pretraining), a novel framework that uses text as a bridge to align unpaired optical and SAR images into a unified embedding space enabling effective text-based retrieval from heterogeneous sources. Our experiments show that CLOSP achieves a new state-of-the-art, improving retrieval nDGC@1000 by 54% over existing models. Additionally, we find that the unified training strategy overcomes the inherent difficulty of interpreting SAR imagery by transferring rich semantic knowledge from the optical domain with indirect interaction. Furthermore, GeoCLOSP, which integrates geographic coordinates into our framework, creates a powerful trade-off between generality and specificity: while the CLOSP excels at general semantic tasks, the GeoCLOSP becomes a specialized expert for retrieving location-dependent crisis events and rare geographic features. This work highlights that the integration of diverse sensor data and geographic context is essential for unlocking the full potential of remote sensing archives.
@@
[…]
@@ 1 Introduction
Satellite remote sensing has become an indispensable tool for monitoring our planet, providing critical data for many applications, including climate science, environmental management, disaster response, urban planning, and precision agriculture (Wulder et al., 2022; Torres et al., 2012; Drusch et al., 2012; Goward et al., 2001; Roy et al., 2014). The large number of satellite missions provides a great volume and diversity of Earth observation data. However, this availability presents a significant challenge: effectively searching and retrieving relevant imagery from massive, heterogeneous, and multimodal archives, particularly when the search criteria are expressed in natural language. To address this challenge, the field of Text-to-Remote-Sensing-Image Retrieval (T2RSIR) has emerged, aiming to bridge the semantic gap between textual descriptions and the rich visual content of satellite imagery. The task involves indexing a vast database of images and retrieving a ranked list of the most pertinent examples that match a user’s natural language query.
The advances in multimodal learning, particularly the adaptation of contrastive learning frameworks like CLIP (Radford et al., 2021), have driven the recent progress in T2RSIR. Models such as SkyCLIP (Wang et al., 2024), SenCLIP (Liu et al., 2024), and RemoteCLIP (Jain et al., 2024) have demonstrated success in aligning text with remote sensing imagery. However, these efforts have focused on high-resolution aerial or RGB satellite images. This reliance on the visible spectrum overlooks the information contained in other modalities and limits their applicability. For instance, missions like Sentinel-1 provide Synthetic Aperture Radar (SAR) provides all-weather, day-and-night acquisitions, making it highly effective for land cover mapping, change detection, and infrastructure monitoring (Clasen et al., 2024; Rege Cambrin and Garza, 2024; Li et al., 2020). Concurrently, multispectral imagery (MSI) from sensors like Sentinel-2 offers crucial spectral bands beyond RGB that are essential for analyzing vegetation health, water bodies, and soil properties (Thenkabail et al., 2004; Gao, 1996; Huete, 1988). While some models like SatCLIP (Klemmer et al., 2024) or Llama3-MS-CLIP (Marimo et al., 2025) have begun to leverage multispectral data, they both disregard the role of SAR data, and SatCLIP aligns it only with geographic coordinates. The full potential of combining these diverse sensor types into a unified retrieval framework remains largely underexplored.
This points to two gaps in the current state-of-the-art. First, existing T2RSIR corpora are inadequate for developing and evaluating rich and robust multimodal systems. Many are limited by small scale (e.g., RSICD, RSITMD (Lu et al., 2018; Yuan et al., 2022)), a dependency on RGB-only aerial imagery, and a lack of rich sensor data like MSI and SAR. Furthermore, their annotations often consist of unstructured, free-form text from non-experts or language models (Wang et al., 2024; Liu et al., 2025), which can introduce ambiguity and limit verifiable quantitative analysis. While datasets with structured Land Use/Land Cover (LULC) annotations exist (e.g., Sen2Lucas (Jain et al., 2024)), they lack crisis-related information and still rely on the optical domain. Second, a direct consequence of this data scarcity is the absence of retrieval models capable of jointly interpreting textual queries, multispectral optical data, and SAR imagery. No existing architecture creates a shared semantic space that unifies these three modalities, preventing the development of retrieval systems that can leverage the unique physical information captured by different sensors.
To bridge this gap, we present CrisisLandMark, a new, large-scale T2RSIR corpus of over 647,000 image-text pairs from the Sentinel-1 (SAR) and Sentinel-2 (optical) missions. Our corpus is enriched with structured, machine-readable annotations for LULC and crisis events, derived from authoritative sources like CORINE Land Cover (Büttner et al., 2017) and Dynamic World (Brown et al., 2022). This provides an unambiguous foundation for training and evaluation. Its characteristics are summarized in Table˜1. Additionally, to leverage this rich dataset, we propose CLOSP (Contrastive Language Optical SAR Pretraining), a novel multimodal architecture designed to align textual descriptions, optical MSI, and SAR data into a shared latent space. By using the structured text annotations as a common anchor, CLOSP effectively learns to associate the different visual modalities without requiring temporally aligned image pairs. Furthermore, we introduce GeoCLOSP, which integrates geographic coordinates into CLOSP to explore the trade-off between semantic generality and geographic specificity.
@@
Our experimental results on the new CrisisLandMark benchmark demonstrate that CLOSP substantially outperforms existing state-of-the-art T2RSIR models. We show that its unified training strategy enables powerful knowledge transfer, significantly improving the semantic retrieval of challenging SAR imagery. The findings underscore the value of integrating diverse sensor data to build more powerful, robust, and versatile retrieval systems for large-scale Earth observation.
In summary, the main contributions of this paper are:
@@
  • A new, large-scale T2RSIR corpus, CrisisLandMark, featuring over 647,000 paired examples of multispectral optical (Sentinel-2) and SAR (Sentinel-1) imagery with structured annotations for land use, land cover, and crisis events.
  • A novel multimodal learning architecture, CLOSP, that is the first to create a shared semantic space aligning textual descriptions, optical MSI, and SAR imagery, with an extension, GeoCLOSP, that incorporates geographic locations.
  • A comprehensive empirical validation demonstrating CLOSP’s superiority over state-of-the-art baselines and quantifying the trade-off between semantic retrieval and geographic specificity.
The code and corpus are publicly available at https://github.com/DarthReca/closp. The remainder of this paper is organized as follows: Section 2 details the construction of the CrisisLandMark corpus. Section 3 presents the CLOSP and GeoCLOSP architectures. Section 4 describes the experimental setup and results, Section 5 discusses the findings, and Section 6 concludes the paper, presenting future works.
@@ 2.1 Data sources
We enrich the source images with a comprehensive set of textual and geospatial annotations. To support crisis management applications, we incorporated the four specialized datasets (CaBuAr, MMFlood, Sen12Flood, Quakeset), retaining their original event tags (wildfire, flood, earthquake). For these crisis images, we generated corresponding LULC annotations (trees, bare, built, crops, grass, water, snow and ice, shrub and scrub, flooded vegetation) by querying the Dynamic World collection. For general LULC coverage in re-BEN, we mapped the high-quality CORINE LC provided with the dataset as detailed in Section˜2.3. Spatial coordinates also characterize all the image patches of every dataset.
The resulting corpus is primarily composed of general land cover scenes from re-BEN, which constitute 88% of the images, providing a robust foundation for learning diverse visual semantics. The remaining 12% from the specialized datasets offer critical examples for crisis retrieval tasks. Table˜2 provides a detailed breakdown of the data sources and their respective contributions of Sentinel-1 and Sentinel-2 imagery.
@@ 2.2 Source Images
Sentinel-2 Level-2A (L2A) products (Drusch et al., 2012; European Space Agency, 2024b) are high-resolution satellite images that provide detailed, true-color views of the Earth’s surface, optimized for environmental and land monitoring. These images are processed to correct atmospheric effects like haze or cloud cover, making them more accurate and ready for analysis. They include 13 spectral bands, ranging from visible light (what we see) to infrared, enabling insights into vegetation health, water quality, soil conditions, and urban development. After the atmospheric correction, one band was removed, resulting in 12 bands, as seen in Table˜3. One example image of the RGB bands can be seen in LABEL:fig:example_general_S2 paired with VV and VH channels (LABEL:fig:example_general_S1_1 and LABEL:fig:example_general_S1_2) of a Sentinel-1 GRD product of the same area. The different vision achieved with different spectral bands is shown in Figure˜3 in the case of a wildfire.
Figure 2: Sample images from Sentinel-1 $V$, and Sentinel-1 VH, and Sentinel-2 RGB of the same area. The scale of Sentinel-1 image values for each channel is the same.
@@ 2.3 Source text
Since CORINE is limited to specific years and only available for Europe, we enrich the four crisis-event datasets (that cover different years and areas than CORINE) with the 9 LULC classes of Dynamic World (DW) (Brown et al., 2022) (i.e., Water, Trees, Grass, Flooded Vegetation, Crops, Shrub and Scrub, Built area, Bare Ground, Snow), which are contained in CORINE categories. The crisis annotation is taken from the original dataset. Dynamic World is a high-resolution, near-real-time global land cover system developed by Google, leveraging deep learning models and Sentinel-2 satellite imagery. It provides 10-meter spatial resolution land cover classifications of the entire globe, making it applicable to diverse ecosystems and regions. Unlike traditional datasets, it offers frequent updates, enabling the detection of rapid and fine-scale land cover changes. The crisis datasets also contain special keywords to denote burned area, earthquake damaged, and flooded area. To create a unified and scalable label space, we harmonized all annotations using the 9-class system from DW. Thanks to its near-real-time global coverage, it is ideal for analyzing diverse ecosystems and detecting rapid land cover changes. With this solution, we map the higher quality of CLC annotation to the same classes of DW. The description and distribution of the classes can be found in Table˜4. The mapping is detailed in Appendix˜C.
@@ 2.4 Queries
We define the query set by taking every unique combination (with length from 1 to 12) of one or more labels that co-occur in at least one image within our corpus. This process resulted in a total of 2,047 distinct multi-label queries. We define a graded relevance relrel based on the IoU between the query labels $L_{q}$ and the image labels $L_{i}$ to provide a more fine-grained evaluation:
@@
For binary relevance operators (e.g., precision and recall), we consider a threshold of 5 for defining a relevant item. The distribution of relevance scores and relevant images per query is shown in Figure˜4.
@@ 3 Methodology
Figure 5: The CLOSP model aligns textual descriptions with SAR and optical satellite images. For a batch of NN elements ($M$ SAR and $M$ MSI), one modality—either SAR or optical—is selected for each element of the batch, and the corresponding image embeddings are paired with their associated textual embeddings. The model is trained to maximize alignment for these positive pairs (represented by white cells in the matrix) while ensuring that negative pairs, formed by combining textual and image embeddings from different items within the batch (represented by grey cells in the matrix), are effectively separated. GeoCLOSP extends the CLOSP architecture by incorporating a location encoder, which aligns the geographical coordinates of an item with the corresponding satellite image, in parallel with the image-text alignment.
The remainder of this section is organized as follows. Section 3.1 formulates the task, whereas Sections 3.2 and 3.3 respectively describe CLOSP and its variant that also leverages geospatial data (GeoCLOSP).